OpenStudy (anonymous):

Form a polynomial, f(x), with real coefficients having the given degree and zeros. Degree: 4; zeros: 4i and 5i Form a polynomial, f(x), with real coefficients having the given degree and zeros. Degree: 4; zeros: 4i and 5i @Mathematics

OpenStudy (anonymous):

(x-4i)(x+4i)(x-5i)(x+5i)=0 (x^2+16)(x^2+25)=0 x^4 + 41x^2 + 400 = 0